Answer items 10 to 12 from the following information:
The mean for skinfold fat (hydrostatically determined) is 25% and the standard deviation is 10%. The coefficient of determination for skinfold fat and hydrostatically determined body fat is .91. Assume that the prediction equation for predicting hydrostatically determined percent fat from skinfolds is percent fat' = skinfolds × 0.5 + 2.

10. Calculate Terry's estimated hydrostatically determined percent fat if Terry’s skinfold thickness is 20 mm.

11. Calculate the error for predicting percent fat from skinfolds.
Answer #1

10)

Terry's estimated hydrostatically determined percent fat =20*0.5+2 =12 %

11)

error for predicting percent fat from skinfolds =sy*√(1-r2) =10*sqrt(1-0.91)=3